OpenStudy (anonymous):

How do I solve this Which of the following is equal to (3x/5-7)+7?

OpenStudy (e.mccormick):

\(\left(\dfrac{3x}{5}-7\right)+7\) ?

OpenStudy (e.mccormick):

If so, associative law of addition: \(\left(\dfrac{3x}{5}-7\right)+7=\dfrac{3x}{5}\left(-7+7\right)\) Now simplify and see if it is closer to any of your answers.

OpenStudy (anonymous):

@Savane0621 did you need help

OpenStudy (anonymous):

Which is an equivalent form of the equation? 3x- 3y = 15

OpenStudy (anonymous):

i think i know how to do this

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=+3x-+3y+%3D+15 look on this site

OpenStudy (e.mccormick):

Why are you tagging someone else into a question that is a year old to ask if they want help with something else?